# Curriculum effects and compositionality emerge with in-context learning in neural networks

### v3

@@ 3.2 Compositionality
Compositionality is thought to be a key property underlying human cognitive flexibility, permitting familiar rules or concepts to be combined in novel ways, thus facilitating a powerful form of generalization [20, 12, 67, 91]. Recent work has shown that although compositionality may not be a natural property of standard IWL in neural networks [20, 39, 69, 40], it can emerge as a property of an ICL algorithm [53, 67]. Our results build on this work, showing that it is possible to endow a neural network with an ICL algorithm that is capable of reproducing the compositional generalization behaviors observed in humans in a recent study [19], even when standard IWL fails (see test accuracy in Figure 3f). We showed that this kind of ICL algorithm can be metalearned by training on a distribution of related tasks, but also emerges in LLMs that are pretrained on large corpora of text (see Figure 4). While metalearning offers a clear understanding of how a neural network can come to implement an emergent compositional learning algorithm [53, 67], it is less clear why this property would emerge in LLMs trained on next-word prediction. One suggestion is that the large-scale pretraining of LLMs can itself be seen as a kind of metalearning [48, 87], where some subset of samples from the training distribution puts pressure on these models to learn how to compose novel concepts presented in context [67]. This is consistent with the hypothesis that human compositionality is metalearned—a conjecture that, while difficult to study, may yield specific empirical predictions [26, 53, 92, 93]. Finally, a key contribution of our work is that it builds on studies linking compositionality to curriculum effects in humans [19], providing a unified account of compositionality, curriculum effects, and their interaction by demonstrating that a single neural network capable of both ICL and IWL can reproduce them.
3.3 Metalearning as a tool for modeling cognition
Metalearning allows a neural network to learn how to learn new tasks through its forward activation dynamics [94, 53, 66, 47]. Increasingly used in cognitive modeling [64], metalearning is a versatile technique that allows the modeler to impart any inductive bias to a neural network so long as it can be implicitly specified by a dataset of input-output pairs [95]. We used metalearning to impart inductive biases for compositionality and the blocking advantage. As others have noted [64, 53, 95], researchers can deploy the metalearning approach to understand the consequences of these inductive biases while remaining agnostic about whether the metalearning process models the evolution of innate biases or their development within an individual’s lifetime. Regardless of their origin in humans, our simulations show how compositionality and curriculum effects can emerge in a single neural network through the interactions between ICL and IWL.
3.4 One network or two systems?
We have emphasized how two competing learning processes can coexist in a single neural network, with one taking place through activation dynamics (ICL), and the other taking place through changes to synaptic weights (IWL). Although the distinction between activation-based and weight-based dynamics is relevant to many aspects of computation in the brain [13, 96, 97, 47], we note that the functional roles we hypothesize for ICL and IWL may also map onto different brain regions. The organization of the human PFC, which has an intrinsic bias to robustly maintain information over longer timescales until it is actively updated [13, 98, 99], may encourage ICL abilities, along with their specific properties, to become localized to this area [100, 36]. Indeed, the PFC is known to be important for the flexible adaptation of behavior to the current context [46], and in the maintenance of task sets [101, 102] and goals [103].
@@ 3.4 One network or two systems?
While we used standard transformers [90] that did not contain any separate PFC-like system, we note that the ICL algorithms implemented in their activation dynamics can be seen as analogous to those observed in neural models of PFC trained across multiple tasks [14, 47, 10]. Recent work has shown that transformer architectures can mimic the frontostriatal gating mechanisms in these biological models when trained on human working memory tasks, and exhibit effective capacity limitations despite the lack of any inherent architectural constraint imposing such a limitation [104, 105]. Future work could use similar techniques to investigate whether emergent PFC-like computational mechanisms also explain ICL-related phenomena in our metalearning networks.
4 Methods
4.1 Task details
@@ 4.2 Model details
In all models, tokens were embedded using a dictionary of learnable vectors. In the metalearning experiments, these embeddings started out as arbitrary random vectors and were optimized by end-to-end backpropagation throughout training. In the LLMs, real English words were used, allowing the models to leverage semantic knowledge gained through pretraining. This is similar to the human participants in the original studies, who could leverage existing knowledge that color and animal are orthogonal feature dimensions, for example.
@@
All metalearning experiments used the same transformer architecture [75, 90]. An informal hyperparameter search was conducted to find a suitable number of layers, hidden size, dropout, and learning rate. The size of the feedforward layers was always twice the hidden size. The best-performing model was selected based on validation accuracy for each task separately. In the category task, the best-performing model had 4 layers, 8 heads, a hidden size of 64, and no dropout. In the compositional task, the best-performing model had 12 layers, 8 heads, a hidden size of 64, and dropout of 0.1. Models were were evaluated on exact-match accuracy using greedy decoding and teacher forcing.
@@
In the LLM experiments, we evaluated GPT-3.5 [48, 76] and Llama 2 [75]. GPT-3.5 is an LLM pretrained on next-token prediction and finetuned to be more useful in a chat-based interface. We used the version of Llama 2 that has not been finetuned on instruction data. In GPT-3.5 (“gpt-3.5-turbo-instruct”), the temperature was set to 0.1, and five runs were performed. A maximum of 7 tokens were generated, and no post-processing was done except to strip extra spaces. Llama 2 is an open-source model with approximately 70 billion parameters. The model was run using resources from the Center for Computation and Visualization at Brown University. The model was quantized so it could fit onto 2 gpus. A number of different prompts for each model were tested, but good performance was achieved with simple prompts containing only the study examples, and the prompts did not qualitatively change the pattern of results across conditions.
4.3 Metalearning
@@
We adopted a metalearning framework to induce an ICL algorithm to emerge within the activation dynamics of a neural network by training it on a distribution of tasks [47, 64, 53]. These task distributions encouraged the resulting ICL algorithm to have a preference for related trials to be blocked over time, and a tendency to generalize compositionally.
@@
For the category-learning experiments, we trained our networks on a distribution of tasks with the same basic structure described above. Each individual task was sampled as follows: 2 feature dimensions were sampled uniformly without replacement from a set of 200 unique dimensions. Each of these dimensions had 8 possible values, making 64 possible items in the newly sampled task. One of two possible category labels was randomly assigned to each of the two categories. In each new task, 16 items from each category were randomly chosen to be included in the set of 32 study examples. The queries seen during metalearning could either be one of the 32 given in the context (“train”), or one of the remaining 32 (“test”). In our main experiments, all samples in the metalearning distribution used the rule-like task and the blocked condition. The network metalearned on 12,000 tasks sampled in this way, and was subsequently tested on a held-out set of 100 tasks with combinations of dimensions that had not been seen during training. A further 10 held-out tasks were used for testing and finetuning. During metalearning in the category setting, networks trained for 20 epochs with the Adam optimizer [106], a learning rate of 0.0001, and a batch size of 256.
@@
We also constructed a metalearning distribution based on the design of the compositional task [19]. Again, each individual task in this distribution had the same structure as the compositional task presented above. The tasks were sampled as follows: First, the orders of the lists of five colors and five animals were shuffled, determining their corresponding orders in the 5x5 grid of locations. Then, the two features were randomly assigned to the x- and the y-coordinates (color = x and animal = y, or vice versa). In the rotated condition, this 5x5 grid was rotated by 45 degrees and scaled so that each coordinate of each cue landed on an integer. As in the category-learning setting, all samples in the metalearning distribution were rule-like and blocked. We again generated 12,000 tasks for metalearning, and used 100 held-out tasks with different 5x5 grids for validation. A further 10 held-out tasks were used for testing and finetuning. During metalearning in the compositional task setting, networks trained for 500 epochs with the Adam optimizer [106], a learning rate of 0.001, and a batch size of 256.
4.4 Finetuning
@@
Once the network acquired an ICL algorithm through metalearning, it was subsequently evaluated on its ability to learn new unseen tasks from each condition. This evaluation was conducted in two ways. In the few-shot evaluation, the weights of the network were frozen, ensuring that all learning was due to ICL on the study examples given in context. In the finetuning evaluation, the model’s weights were not frozen, and any errors made were backpropagated to update weights. During finetuning, the model only received feedback on the study examples, thus emulating the experience of the human participants [19]. Note that this is unlike the metalearning phase, when the model learned how to generalize to queries not included in the study examples. This second IWL learning phase that the model underwent can be understood as ‘finetuning’ the model on a specific task, while the metalearning can be understood as ‘pretraining.’
@@
During the finetuning phase, samples were either blocked or interleaved in two distinct but congruent ways. In the blocked condition, related items were blocked over the context, but they were also blocked over the gradient steps (i.e., the model was finetuned for N gradient steps on samples containing queries from one stimulus group, then was finetuned for N gradient steps on samples containing queries from the other group, and so on). Likewise, in the interleaving condition, items from each group were interleaved both over the context and over the gradient steps. In the main experiments, the curriculum condition was always consistent during finetuning—related items were either blocked over both the context and the gradients steps, or interleaved over both the context and the gradient steps. However, for the sake of completeness we experimented with all combinations and report these results in the Appendix.
@@ A.2 Text-based versions of the tasks
A.3 Metalearning details
Metalearning networks were trained on distributions of tasks and evaluated on tasks that were not seen during training. In both the metalearning and the finetuning phases, networks received text-based inputs like the ones described above and predicted the outputs corresponding to the test query (i.e., the 1-token category label following the test query in the category-learning task, or the 2-token reward location following the test query in the compositional task). Cross-entropy loss was computed at only these specific output tokens.
In both task settings, 10 different runs of metalearning were performed with different random initializations. Unless otherwise specified (see Additional Experiments below), metalearning was always done with task distributions that had rule-like structure and blocked curricula. After metalearning was complete, models were evaluated in the few-shot setting on 100 validation tasks and were finetuned on 10 test tasks from each of the four conditions (Rule-like and Blocked, Rule-like and Interleaved, Rotated and Blocked, Rotated and Interleaved). Particular seeds were excluded if they did not pass a 90% accuracy threshold in the few-shot evaluation, because our main questions of interest pertained to models that had acquired a competence in learning in-context. In the IWL-only experiments (see Figure 2d-f and Figure 3d-f in the main text), the metalearning step was skipped, and 10 different randomly initialized networks were finetuned on 10 test tasks from each condition in the same way.
@@ B.1 IWL exhibits an interleaving advantage on the category task
We hypothesized that simple in-weight learning on the category-learning task would show an interleaving advantage, regardless of rotation. To test this hypothesis, we trained networks from scratch in each of the four conditions (Rule-like and Blocked, Rule-like and Interleaved, Rotated and Blocked, Rotated and Interleaved) using the same setup used for finetuning networks after metalearning. We analyzed the results using a generalized linear model (GLM) with a quasi-binomial family using the logit link function. This model predicted the train accuracy (successes vs. failures) across the four experimental conditions from the Rotation condition (Rule-like vs. Rotated), the Curriculum condition (Blocked vs. Interleaved), and their interaction (see Figure 2e in the main text). The results revealed a significant main effect of Curriculum ($\chi^{2}=786.88$, $p<2.2e-16$) and a significant interaction between Rotation and Curriculum ($\chi^{2}=31.18$, $p=2.349e-08$). The main effect of Rotation was not significant ($\chi^{2}=0.0$, $p=1.0$). The significant interaction was due to a small but statistically significant difference in the final interleaved accuracy between the two rotation conditions, indicating that 100% accuracy was slightly harder to achieve on the rotated version of the task (see Table 1). We therefore tested the simple main effects of Curriculum in each of the Rotation conditions, and found a significant interleaving advantage in both the Rule-like ($\chi^{2}=7e+16$, $p<2.2e-16$) and Rotated ($\chi^{2}=393$, $p<2.2e-16$) conditions. Overall, these findings confirm our hypothesis that in the absence of in-context learning, in-weight learning shows an interleaving advantage that is largely robust to the rotation condition.

### v4

@@ Discussion
Compositionality is thought to be a key property underlying human cognitive flexibility, permitting familiar rules or concepts to be combined in novel ways, thus facilitating a powerful form of generalization [21, 12, 70, 107]. Recent work has shown that although compositionality may not be a natural property of standard IWL in neural networks [21, 37, 72, 38], it can emerge as a property of an ICL algorithm [55, 70]. Our results build on this work, showing that it is possible to endow a neural network with an ICL algorithm that is capable of reproducing the compositional generalization behaviors observed in humans in a recent study [20], even when standard IWL fails (see test accuracy in Figure 3f). We showed that this kind of ICL algorithm can be metalearned by training on a distribution of related tasks, but also emerges in LLMs trained on large corpora of text (see Figure 4). While metalearning offers a clear understanding of how a neural network can come to implement an emergent compositional learning algorithm [55, 70], it is less clear why this property would emerge in LLMs trained on next-word prediction. One suggestion is that at large enough scales, the language modeling objective used in LLMs can itself be seen as engendering a kind of metalearning [50, 102], where some subset of training samples puts pressure on these models to learn how to compose novel concepts or reasoning steps in context [70]. This is consistent with the hypothesis that human compositionality is metalearned—a conjecture that, while difficult to study, may yield specific empirical predictions [27, 55, 108, 109]. Finally, a key contribution of our work is that it builds on studies linking compositionality to curriculum effects in humans [20], providing a unified account of compositional generalization and its dependence on curriculum.
One network or two systems for category learning?
While our neural networks are not meant to be comprehensive models of human category learning, they may be relevant to other phenomena observed in category-learning studies. One ongoing debate in this area has been about whether human category learning is best characterized by a single learning system or by multiple systems [110, 87, 1, 111, 112, 113, 114, 115, 116, 117]. Single-system theories emphasize the principle of parsimony, and argue that a system that relies on stimulus similarity and selective attention can explain most of the available findings [112]. Multiple-systems theories argue that a single system is not sufficient to account for double dissociations evident in human behavior [1, 118, 119], such as the one pertaining to the curriculum effects discussed above [5].
Our work may help resolve this debate by showing how such double dissociations can be explained by a single network that can learn in two different ways. ICL and IWL are not separate learning systems, but nevertheless manifest fundamentally different properties and compete to drive learning behavior, with each taking precedence at different times. Our approach arguably maintains the parsimony of a single-system theory in the sense that these two distinct sets of learning properties emerge from the natural dynamics of a single network, rather than being independently posited as part of separate systems. However, as discussed above, the properties of ICL and IWL align well with the two systems proposed in prominent multiple-system theories [110, 1], with ICL corresponding to the explicit, verbal system and IWL corresponding to the implicit, procedural system.
In addition to the curriculum effects we observed in our experiments, the distinction between ICL and IWL may help to explain other findings motivating multiple-system theories of category learning. For example, some studies have shown that increased WM load can impair rule-based learning [118, 120, 121, 122, although see 123, 114]. This finding parallels our results showing that ICL-mediated generalization suffers when access to contextual information is restricted (see Figure 5). Studies of category learning have also shown that humans are capable of generalizing outside of their training distribution on rule-based tasks, but cannot do so in information-integration (rotated) tasks [124, 125]. This aligns with our finding that ICL, but not IWL, was capable of generalizing on both tasks (see Figure 3). Finally, there is some evidence that children struggle specifically with rule-based category-learning tasks [126, 127, 128, 129], but can perform at adult levels when categories are based on family resemblances [130]. This is consistent with our neural networks, which are inherently capable of IWL but only develop sophisticated ICL through metalearning [27].
Our models may also clarify certain outstanding questions for current multiple-system theories. For example, behavioral evidence suggests that the verbal or explicit system operates by default initially in humans, but it is unclear a priori why this would be the case [5, 1]. In our neural network models, ICL operates by default because it can occur at a much faster timescale (through activation dynamics), and because IWL only occurs when errors are made. Another unresolved question concerns evidence from neuroimaging studies on category learning suggesting that there is substantial overlap in the brain regions active during rule-based and information-integration tasks [131, 132]. This can seem to contradict the predictions of a multiple-system theory that posits completely independent learning modules. The distinction between ICL and IWL provides a natural explanation for this finding, as these two learning processes coexist throughout the network and therefore need not be localizable to separate regions.
In fact, our neural networks are likely to be unrealistically homogeneous, as they have no inherent modularity at all. Many findings suggest that specific brain regions such as PFC are particularly important for cognitive functions such as WM, rule-based inference, and modulating processing according to the current context or goal [133, 48, 134, 14, 135]. We speculate that the organization of the human PFC, which has an intrinsic bias to robustly maintain information over longer timescales until it is actively updated [13, 136, 137], may encourage ICL abilities, along with their specific properties, to become partially localized to this area [138, 42].
@@
Although our models did not contain any separate PFC-like system, we note that the ICL algorithms implemented in their activation dynamics can be seen as analogous to those observed in neural models of PFC trained across multiple tasks [14, 49, 10]. Just as in our models, these ICL-like abilities only emerge through IWL-like learning of abstract representations in PFC and of gating policies in BG. Recent work has shown that transformer architectures can mimic the frontostriatal gating mechanisms in these biological models when trained on human WM tasks, and exhibit effective capacity limitations despite the lack of any inherent architectural constraint imposing such a limitation [139, 140]. Future work could use similar techniques to investigate whether emergent PFC-like computational mechanisms also explain ICL-related phenomena in our metalearning networks.
Methods
Task details
@@ Methods
In all models, tokens were embedded using a dictionary of learnable vectors. In the metalearning experiments, these embeddings started out as arbitrary random vectors and were optimized by end-to-end backpropagation throughout training. In the LLMs, real English words were used, allowing the models to leverage semantic knowledge gained through training. This is similar to the human participants in the original studies, who could leverage existing knowledge that color and animal are orthogonal feature dimensions, for example.
@@
All metalearning experiments used the same transformer architecture [78, 105]. An informal hyperparameter search was conducted to find a suitable number of layers, hidden size, dropout, and learning rate. The size of the feedforward layers was always twice the hidden size. The best-performing model was selected based on validation accuracy for each task separately. In the category task, the best-performing model had 4 layers, 8 heads, a hidden size of 64, and no dropout. In the compositional task, the best-performing model had 12 layers, 8 heads, a hidden size of 64, and dropout of 0.1. Models were evaluated on exact-match accuracy using greedy decoding and teacher forcing.
@@
In the LLM experiments, we evaluated GPT-3.5 [50, 79] and Llama 2 [78]. GPT-3.5 is an LLM trained on next-token prediction and finetuned to be more useful in a chat-based interface. We used the version of Llama 2 that has not been finetuned on instruction data. In GPT-3.5 (“gpt-3.5-turbo-instruct”), the temperature was set to 0.1, and five runs were performed. A maximum of 7 tokens were generated, and no post-processing was done except to strip extra spaces. Llama 2 is an open-source model with approximately 70 billion parameters. The model was run using resources from the Center for Computation and Visualization at Brown University. The model was quantized so it could fit onto 2 gpus. A number of different prompts for each model were tested, but good performance was achieved with simple prompts containing only the study examples, and the prompts did not qualitatively change the pattern of results across conditions.
Metalearning
@@
We adopted a metalearning framework to induce ICL abilities to emerge within the activation dynamics of a neural network by training it on a distribution of tasks [49, 66, 68, 55]. These task distributions encouraged the resulting ICL algorithm to have a preference for related trials to be blocked over time, and a tendency to generalize compositionally.
@@
For the category-learning experiments, we trained our networks on a distribution of tasks with the same basic structure described above. Each individual task was sampled as follows: 2 feature dimensions were sampled uniformly without replacement from a set of 200 unique dimensions. Each of these dimensions had 8 possible values, making 64 possible items in the newly sampled task. One of two possible category labels was randomly assigned to each of the two categories. In each new task, 16 items from each category were randomly chosen to be included in the set of 32 study examples. The queries seen during metalearning could either be one of the 32 given in the context (“train”), or one of the remaining 32 (“test”). In our main experiments, all samples in the metalearning distribution used the rule-like task and the blocked condition. The network metalearned on 12,000 tasks sampled in this way, and was subsequently tested on a held-out set of 100 tasks with combinations of dimensions that had not been seen during training. A further 10 held-out tasks were used for testing. During metalearning in the category setting, networks trained for 20 epochs with cross-entropy loss, the Adam optimizer [141], a learning rate of 0.0001, and a batch size of 256.
@@
We also constructed a metalearning distribution based on the design of the compositional task [20]. Again, each individual task in this distribution had the same structure as the compositional task presented above. The tasks were sampled as follows: First, the orders of the lists of five colors and five animals were shuffled, determining their corresponding orders in the 5x5 grid of locations. Then, the two features were randomly assigned to the x- and the y-coordinates (color = x and animal = y, or vice versa). In the rotated condition, this 5x5 grid was rotated by 45 degrees and scaled so that each coordinate of each cue landed on an integer. As in the category-learning setting, all samples in the metalearning distribution were rule-like and blocked. We again generated 12,000 tasks for metalearning, and used 100 held-out tasks with different 5x5 grids for validation. A further 10 held-out tasks were used for testing. During metalearning in the compositional task setting, networks trained for 500 epochs with the Adam optimizer [141], a learning rate of 0.001, and a batch size of 256.
Task-specific training
@@
Once the network acquired an ICL algorithm through metalearning, it was subsequently evaluated on its ability to learn new unseen tasks from each condition. This evaluation was conducted in two ways. In the few-shot evaluation, the weights of the network were frozen, ensuring that all learning was due to ICL on the study examples given in context. In task-specific training, the model’s weights were not frozen, and any errors made were used to update weights. During task-specific training, the model learned a single task and only received feedback on the study examples, thus emulating the experience of the human participants [20]. Note that this is unlike the metalearning phase, when the model learned how to generalize to queries not included in the study examples. This second task-specific learning phase that the model underwent can be understood as ‘finetuning’ the model on a specific task, while the metalearning can be understood as ‘pretraining.’ During task-specific training, networks were again trained with cross-entropy loss and the Adam optimizer [141], with a learning rate of 0.00001 in the category-learning task, and a learning rate of 0.0001 in the compositional task. In both tasks, the batch size was equal to the total number of examples (i.e., queries) used in a given block (32 in the category-learning setting, 5 in the compositional setting).
@@
During the task-specific training phase, samples were either blocked or interleaved in two distinct but congruent ways. In the blocked condition, related items were blocked over the context, but they were also blocked over the gradient steps (i.e., the model was trained for N gradient steps on samples containing queries from one stimulus group, then was trained for N gradient steps on samples containing queries from the other group, and so on). Likewise, in the interleaving condition, items from each group were interleaved both over the context and over the gradient steps. In the main experiments, the curriculum condition was always consistent during task-specific training—related items were either blocked over both the context and the gradients steps, or interleaved over both the context and the gradient steps. However, for the sake of completeness we experimented with all combinations and report these results in the Appendix.
@@ A.2 Text-based versions of the tasks
A.3 Training details
Metalearning networks were trained on distributions of tasks and evaluated on tasks that were not seen during training. In both the metalearning and the task-specific training phases, networks received text-based inputs like the ones described above and predicted the outputs corresponding to the test query (i.e., the 1-token category label following the test query in the category-learning task, or the 2-token reward location following the test query in the compositional task). Cross-entropy loss was computed at only these specific output tokens. During metalearning, models were trained with the Adam optimizer
In both task settings, 10 different runs of metalearning were performed with different random initializations. Unless otherwise specified (see Additional Experiments below), metalearning was always done with task distributions that had rule-like structure and blocked curricula. After metalearning was complete, models were evaluated in the few-shot setting on 100 validation tasks and independently underwent task-specific training (finetuning) on 10 different test tasks from each of the four conditions (Rule-like and Blocked, Rule-like and Interleaved, Rotated and Blocked, Rotated and Interleaved). During task-specific training, Particular seeds were excluded if they did not pass a 90% accuracy threshold in the few-shot evaluation, because our main questions of interest pertained to models that had acquired a competence for in-context learning. In the IWL-only experiments (see Figure 2d-f and Figure 3d-f in the main text), the metalearning step was skipped, and 10 different randomly initialized networks underwent task-specific training on 10 separate test tasks from each condition in the same way.
A.4 Tradeoff between flexibility and retention
In the experiments studying the emergent tradeoff between flexibility and retention (Figure 5 in the main text), we restricted the amount of contextual information available to the model by “ablating” attention to the examples available in the context. This was achieved by masking attention to each of these examples independently with probability $p_{a}$. Masking tokens in this way sets the attention distributions over them to 0, thus disallowing representations encoding subsequent positions in the sequence from extracting any information from them. In our experimental setup, this was analogous to increasing working memory load or reducing working memory capacity, because when attention to particular examples is masked, they can no longer be used to inform the models’ predictions for the query. We also tried applying noise to the attention mechanism (the value vectors) itself, which produced qualitatively similar results (see Additional results section below). Attention was ablated at different levels of $p_{a}$ throughout the task-specific training of models that had already undergone metalearning on unrotated tasks with interleaved curricula. In the flexibility test (leftmost plot in Figure 5 of main text), accuracy on held-out examples was tested under those same values for $p_{a}$. In the retention test (rightmost plot in Figure 5 of the main text), models were evaluated on seen examples without any contextual information available ($p_{a}=1.0$). This simulated a test after a long delay, where no information stored in working memory could be leveraged to make predictions. We tested this tradeoff in both the category-learning task and the compositional task, but due to space limitations, we only report the results for the compositional task in the main text. The results for the category-learning task are shown below (see Additional results section).
@@ B.1 IWL exhibits an interleaving advantage on the category task
We hypothesized that simple in-weight learning on the category-learning task would show an interleaving advantage, regardless of rotation. To test this hypothesis, we trained networks from scratch in each of the four conditions (Rule-like and Blocked, Rule-like and Interleaved, Rotated and Blocked, Rotated and Interleaved) using the same setup used for task-specific training following metalearning. We analyzed the results using a generalized linear model (GLM) with a quasi-binomial family using the logit link function (no random effects). This model predicted the train accuracy (successes vs. failures) across the four experimental conditions from the Rotation condition (Rule-like vs. Rotated), the Curriculum condition (Blocked vs. Interleaved), and their interaction (see Figure 2e in the main text). The results revealed a significant main effect of Curriculum ($\chi^{2}=786.88$, $p<2.2e-16$) and a significant interaction between Rotation and Curriculum ($\chi^{2}=31.18$, $p=2.349e-08$). The main effect of Rotation was not significant ($\chi^{2}=0.0$, $p=1.0$). The significant interaction was due to a small but statistically significant difference in the final interleaved accuracy between the two rotation conditions, indicating that 100% accuracy was slightly harder to achieve on the rotated version of the task (see Table 1). We therefore tested the simple main effects of Curriculum in each of the Rotation conditions, and found a significant interleaving advantage in both the Rule-like ($\chi^{2}=7e+16$, $p<2.2e-16$) and Rotated ($\chi^{2}=393$, $p<2.2e-16$) conditions. Overall, these findings confirm our hypothesis that in the absence of in-context learning, in-weight learning shows an interleaving advantage that is largely robust to the rotation condition.